OpenStudy (anonymous):

how do i integrate cos^2x ?

OpenStudy (anonymous):

standard trick is to rewrite using one of those trig reduction formulas that you thought you would never use

OpenStudy (anonymous):

\[\cos^2(x)=\frac{1}{2}(\cos(2x)+1)\] integrate term by term

OpenStudy (anonymous):

would the answer be 1/20 [1/2 sin 2x ] ?

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

\[\int 1dx=x\] \[\int \cos(2x)dx=-\frac{1}{2}\sin(2x)\]

OpenStudy (anonymous):

sorry last one is wrong should be \[\int \cos(2x)dx=\frac{1}{2}\sin(2x)\]

OpenStudy (anonymous):

ohh, i took out the one outside.

OpenStudy (anonymous):

should get as a final answer \[\frac{x}{2}+\frac{1}{2}\sin(2x)\]

OpenStudy (anonymous):

i gt 1/40 sin 2x + 1/20x

OpenStudy (anonymous):

or if you like you can rewrite as \[\frac{x}{2}+\frac{1}{2}\sin(x)\cos(x)\] by that double angle formula. either way

OpenStudy (anonymous):

damn another mistake, hold on

OpenStudy (anonymous):

\[\frac{x}{2}+\frac{1}{4}\sin(2x)\] is correct

OpenStudy (anonymous):

i assume the zeros in your answer are typos.

OpenStudy (anonymous):

so the answer would be pi/2?

OpenStudy (anonymous):

the answer is a function, or rather a class of funcitons, not a number

OpenStudy (anonymous):

sorry i have to substitute pi and 0, as the last part.

OpenStudy (anonymous):

thanks a lot for your help !